OpenStudy (mtwizzle420):

what is the side length b in the right triangle below? @roz_mar23

OpenStudy (mtwizzle420):



Nnesha (nnesha):

because there is right triangle so that's why Pythagorean theorem, \[\huge\rm a^2 + b^2 = c^2\] use this formula to find side length b

Nnesha (nnesha):

and remember c is the longest side of right triangle which is hypotenuse

OpenStudy (mtwizzle420):

So @Nnesha I got 7^2 = 49 12^2= 144 So 49+b^2= 144

Nnesha (nnesha):

yes right now subtract 49 both side

OpenStudy (mtwizzle420):

144-49=95

Nnesha (nnesha):

yes right now take square root both side